OpenStudy (anonymous):

A right circular cylinder has a base whose area is 154 square feet. If the height of the cylinder is 19 feet, what is its lateral surface area? Use π = 22/7

OpenStudy (anonymous):

836 ft^2 990 ft^2 1672 ft^2 1980 ft^2

OpenStudy (anonymous):

Can you also help me with this one please? Hena stores coffee powder in a steel container with a capacity of 82.5 cubic inches. He uses a glass container to store sugar. The two containers are congruent cylinders. If the glass container has a height of 4.2 inches, what is its radius? 2.5 inches 4.2 inches 4.4 inches 8.4 inches

OpenStudy (anonymous):

We know that the area of a circle is given by pi * radius squared, so we can work out the radius of the cylinder as \[\sqrt{area/\pi}\]. We can then work out the circumference of the circle as this is given by\[ 2\pi r\]. We can substitute our expression for the radius into this to give:\[Circumference = 2\pi \sqrt{area/\pi}\] which simpifies to:\[2\sqrt{154\pi}\]Substituting our value for the area and for pi we get \[2\times\sqrt{154\times22/7} \] Since 154 is 22 * 7 we get 22 for the root and a total circumference of 44. The lateral area is then this times the height of the cylinder (imagine the cylinder unrolled into a rectangle) which gives 44 * 19 or 836.